# Randomised benchmarking for universal qudit gates

### v1

@@ C.3 Identification of the irreps
The standard representation of $S_{d}$ is defined by its action on a basis. The action of this representation is permuting the entries of a vector; the permutation applied is $\sigma\in S_{d}$. Because there is a non-trivial invariant subspace of the standard representation, namely the vector space corresponding to those vectors with each entry equal, it is reducible. However, the representation obtained by restricting to the orthogonal complement of $\operatorname{span}\{(1,\dots,1)\}$, denoted by $\operatorname{span}\{(1,\dots,1)\}^{\bot}$, is irreducible—this irrep is called the standard irrep of $S_{d}$.

### v2

@@ C.3 Identification of the irreps
The standard representation of $S_{d}$ is defined by its action on a basis. The action of this representation is by permuting the entries of a vector; the permutation applied is $\sigma\in S_{d}$. Because there is a non-trivial invariant subspace of the standard representation, namely the vector space corresponding to those vectors with each entry equal, it is reducible. However, the representation obtained by restricting to the orthogonal complement of $\operatorname{span}\{(1,\dots,1)\}$, denoted by $\operatorname{span}\{(1,\dots,1)\}^{\bot}$, is irreducible—this irrep is called the standard irrep of $S_{d}$.